Grade 11Physical Chemistry

Find rms velocity of
1)CO2 gas at 27°c
2)O2 gas at 27°c
Answer with proof

MARTIN ABRAHAM
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Kinetic Energy (Translational, in all directions) of a gas as per Kinetic theory of gases at abs. Temperature T:
          per molecule:  KE = 3/2 * k_B T
             per mole:      KE = 3/2 * R * T
            k_B = Boltzmann's constant. R= Universal gas constant.
 Kinetic Energy of a gas (as per mechanics)
                    KE = 1/2 m (Vrms)²
               or  KE = 1/2 M (Vrms)²
       m = mass of a molecule, M = Molar mass of the gas.
Hence,  (Vrms)² = 3 k_B T /m = 3 R T /M
R = 8.3 J/°°K/mole 
M = 44 gm/mole  for CO2  = 0.044 kg/mole
    (Vrms)² = 3 * 8.314 J/°K/mole * (273+27)° / (0.044 kg/mole)
        Vrms = 412.38 m/sec